OpenStudy (anonymous):

Decompose into partial fractions. (4x+34)/(x^2 - 5x - 24)

OpenStudy (nikvist):

\[\frac{4x+34}{x^2-5x-24}=\frac{4x+34}{(x-8)(x+3)}=\frac{A}{x-8}+\frac{B}{x+3}\] \[4x+34=A(x+3)+B(x-8)=(A+B)x+3A-8B\] \[\Rightarrow\quad A+B=4\,\,,\,\,3A-8B=34\quad\Rightarrow\quad A=6,B=-2\] \[\frac{4x+34}{x^2-5x-24}=\frac{6}{x-8}+\frac{-2}{x+3}\]

OpenStudy (anonymous):

Thanks nikvist...can u help me w/ another problem????

OpenStudy (anonymous):

wait im not sure how you got a = 6 adn b = -2

OpenStudy (anonymous):

nvr mind

OpenStudy (nikvist):

where is another problem? :-)

OpenStudy (anonymous):

Expand and simplify: \[\sum_{n=0}^{4} n ^{2}/2\]

OpenStudy (nikvist):

it is simple \[\sum\limits_{n=0}^{4}\frac{n^2}{2}=\frac{1}{2}\sum\limits_{n=0}^{4}n^2=\frac{1}{2}(0^2+1^2+2^2+3^2+4^2)=\frac{1}{2}\cdot 30=15\]

OpenStudy (anonymous):

i have another one

OpenStudy (anonymous):

\[\sum_{n=1}^{3} 1/n!\]

OpenStudy (anonymous):

im confused about teh factorial

OpenStudy (nikvist):

\[\frac{1}{1!}+\frac{1}{2!}+\frac{1}{3!}=\frac{1}{1}+\frac{1}{2\cdot 1}+\frac{1}{3\cdot 2\cdot 1}=1+\frac{1}{2}+\frac{1}{6}=\frac{5}{3}\]

OpenStudy (anonymous):

About factorials. How would you simplify this expresion: 3(n+1)! / 5n!

OpenStudy (anonymous):

do u distribute